Chelsea left the White House and traveled towards the capital at an average speed of 34 km/h. Jasmine left at the same time and traveled in the opposite direction with an avg speed of 65 km/h. Find the number of hours Jasmine needs to travel before they are 59.4 km apart.
I got T = .91 hours Is this correct?

Found 4 solutions by checkley71, edjones, ionesco, bucky:
Answer by checkley71(8403)

34T+65T=59.4
99T=59.4
T=59.4/99
T=.6 HOURS THEY WILL BE 59.4 KM APART.
PROOF
34*.6+65*.6=59.4
20.4+39=59.4
59.4=59.4

Answer by edjones(7311)

No.
Since they are going in opposite directions you add the speeds.
d=tv distance =time * velocity
d=59.4 v=34+65=99
59.4=99t
divide 99 into both sides. t=59.4/99=.6 hrs
.6 hrs=.6*60=36 minutes
EdJones

Answer by ionesco(12)

I hope this is the answer:
d=59.4
d=d1+d2 where d1=v1*t=34*t and d2=v2*t=65*t
then 59.4=34*t + 65*t then t=59.4/(34+65)

Answer by bucky(2100)

Let's see if your answer is correct. In 0.91 hours how far does Chelsea go if she is driving
at 34 km/hr. Use the formula D = r*T where r is 34 km/hr and T is 0.91. The distance
Chelsea travels is D = r*T = 34*0.91 = 30.94 km.
.
Meanwhile Jasmine is going in the opposite direction at a rate of 65 km/hr. So in 0.91 hrs
she goes D = r*T = 65*0.91 = 59.15 km.
.
So in 0.91 hours the distance between the two drivers is the sum of the distances that each
has traveled. They are separated by 30.94 + 59.15 = 90.09 km. Something appears to be wrong
with your solution because they are separated by 90.09 km, and they are only supposed
to be separated by 59.4 km.
.
Let's work the problem.
.
Since we are trying to find the time, let's use the distance equation for each of the
two drivers.
.
In T hours the distance that Chelsea goes is given by D = 34 * T
.
And in the same T hours the distance that Jasmine goes is given by D = 65*T
.
But if we add these two distances they are supposed to total 59.4 km. So if we add the
two right sides of these equations for distance the sum should equal 59.4 km.
.
When we add the two right sides and set the sum equal to 59.4 we get:
.
34*T + 65*T = 59.4
.
If we add the two terms on the left side we get 34T + 65T = 99T and the equation then becomes:
.
99T = 59.4
.
Now we can solve for T by dividing both sides by 99 to get:
.
T = 59.4/99 = 0.6 hours
.
We have an answer of T = 0.6 hours
.
Let's check it.
.
In 0.6 hours at 34 km per hour Chelsea goes 34*0.6 = 20.4 km
.
And in 0.6 hours at 65 km per hour Jasmine goes 65*0.6 = 39.0 km
.
If you add these two distances together you get a total of 59.4 km, just like the problem
said it should be.
.
Hope this helps you to understand the problem a little better and that you are able to
find your error.